OpenStudy (anonymous):

factoring a trinomial of the form ax2+bx+c where a and b are both positive and c is negative you need to find two numbers m and n such that ax2+b+c=ax2+mx+nx+c and

OpenStudy (anonymous):

i doing think thats the full question/

OpenStudy (anonymous):

What's your question @taylor0214 ?

OpenStudy (whpalmer4):

and \(m*n = a*c\)

OpenStudy (whpalmer4):

and \(m+n = b\)